OpenStudy (anonymous):

Let A) 1 -3 -5 1 1 -2 1 -3 1 1 1 4 and b= -6 1 1 6

OpenStudy (anonymous):

a) apply the Gram-Schmidt process to determine an orthonormal basis for R(A)

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

so b is the basis for the solution space

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

so is b only one vector?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i am not sure, but i think we can apply an alternative form of the gram-schmidt, but if anybody else has any ideas, feel free to chime in

OpenStudy (amistre64):

last time i did something like this i crossed v1xv2 and v1xv3 to find othogonal vectors dunno if itll work here tho;

OpenStudy (amistre64):

dunno how we would cross 2, 4component tho .. yet :)

OpenStudy (anonymous):

from what i have seen with the gram shmidth , B has two vectors at least

OpenStudy (amistre64):

matrix A is already linear independant; it just aint orthogonal

OpenStudy (amistre64):

had to look up what an orthonormal was; its an orthogonal set of vectors that are unit length

OpenStudy (amistre64):

what is vector b? is it defined in terms of the given matrix, or is it defined in termsof the orthonromal components?

OpenStudy (amistre64):

or is b in terms of the standard components of 1000, 0100, 0010

OpenStudy (anonymous):

I am reading up on chapter too

OpenStudy (amistre64):

the GS v1 = x1; W1 = span{x1} v2 = x2 - proj(W1)x2 -3 x2.x1 1 - ----- x1 -3 x1.x1 1

OpenStudy (amistre64):

1 1 1 1 -3 1 -3 1 ----------- -3+1-3+1 = -4 ---------- --- = -1 1+1+1+1 = 4 v2 = -3 1 -2 1 1 2 -3 + 1 = -2 1 1 2

OpenStudy (amistre64):

v3 = x3 - proj(W2) x3 which gets tricky

OpenStudy (amistre64):

W2 = span{v1,v2}

OpenStudy (amistre64):

proj(W2)x3 = the sum of the same process above but using x3 in place of x2, and dotting it to v1 and then v2

OpenStudy (amistre64):

-5-4 1 4 1 1 1 1 --------- -5-4+1+4 = -4 ---------- --- <1,1,1,1> = <-1,-1,-1,-1> 1+1+1+1 = 4 -5-4 1 4 -2 2-2 2 --------- -10-8-2+8 = -12 ----------- --- <-2,2,-2,2> = <3,-3,3,-3> 4+4+4+4 = 16 -1-1-1 -1 3 -3 3 -3 --------- <2,-4,2,-4> = v3

OpenStudy (amistre64):

\[\begin{vmatrix}1&-2&2\\1&2&-4\\1&-2&2\\1&-2&-4 \end{vmatrix}\] would be the orthogonal basis; the orthonormal is to unit length those

OpenStudy (amistre64):

middle bottom; not negative; just 2

OpenStudy (anonymous):

thanks , amistre I will study this

OpenStudy (amistre64):

i think i made a mistake in the v3 parts

OpenStudy (amistre64):

what i have for v3 is spose to be subtracted from x3 i think

OpenStudy (amistre64):

-5-4 1 4 x3 -2,4,-2,4 -(badv3) ---------- -7,0,-1,8 = v3

OpenStudy (amistre64):

again, these are orthogonal basis; orthoNormal is same but unit these vectors